OWL AUG-12-99 shift summary

Alexander Ledovskoy (ledovsko@fnpxsr)
Thu, 12 Aug 1999 07:12:48 -0500

Shift Summary for OWL shift, 12-AUG-99

On Shift: Earl, Sasha L.

Hours of beam available at all intensities: 7.00h
Hours of beam available at nominal intensity: 7.00h
Hours of E832 data (special runs not included): 5.83h
Hours of special runs (laser scans, muon runs...): 0.00h

Accounting of lost time: fresh start, changing tapes, resetting DDD4

========================================================================
Runs ended this shift:
R14199 - 22.8 M events Beam_832
========================================================================
Tapes brought to FCC this shift: none
========================================================================
Accesses during shift: none
========================================================================
Overview of shift: E832 data taking until DDD4 problems appear
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
1419 294 0:00 Taking over smooth running

12:05 Online histos:

o DC planes 9 and 4 hit multiplicities are
NOT happy
o HCC threshold has unhappy face but it looks OK

313 CsI 1913 RMS is 33.3% Next spill is OK.

341 Ktevana errors. KFREND. Problem fixed.

347 SEM=1.2E13 Call MCR. The SEM measurement is wrong
(obviously). This happens from time to time, they said.

375 L2BEAM scaler = 0.0 for this spill

401 LIVETIME=0.47 at SEM=6.3E+12 this spill

3:15 Shift check list is done. Absorber NM2ABS2 reads 0
(set to 1). In fact, it does not update it's reading.
The smooth running indicates that we do have the
absorber in place.

5:10 Csi 166 has comparator problems:
Freq=6, range=2.00, cap=2.00

5:20 Check CsI 166 again:
Freq=7, range=2.00, cap=2.00

554 5:50 Triggers stopped: pipeline cr 4 FULL + DDD VETO 4
Tapes are almost full. Shortly after that, the beam
stop went down spontaineously. We took it as a sign to
stop the run, change the tapes and do a fresh start.

Fresh start hangs. Call Ed. ktev4 plane 0 DC has no
heartbeat. Reset from the back. Hearbeat is kicking.

Fresh start does not work again. Power cycle DDD4 from
the front panel. Fresh start. Success.

14202 7:00 Started.

========================================================================
Follow up Notes on Previous Shift:
========================================================================
Items to Follow up on Next Shift:

o CsI 166 : look for comparator problems